The following selected transactions relate to liabilities of United Insulation Corporation. United’s fiscal year ends on December 31.

2018

Jan.

13

Negotiated a revolving credit agreement with Parish Bank that can be renewed annually upon bank approval. The amount available under the line of credit is $30.0 million at the bank’s prime rate.

Feb.

1

Arranged a three-month bank loan of $8.4 million with Parish Bank under the line of credit agreement. Interest at the prime rate of 10% was payable at maturity.

May

1

Paid the 10% note at maturity.

Dec.

1

Supported by the credit line, issued $18.3 million of commercial paper on a nine-month note. Interest was discounted at issuance at a 9% discount rate.

31

Recorded any necessary adjusting entry(s).

2019

Sept.

1

Paid the commercial paper at maturity.


Required:
Prepare the appropriate journal entries through the maturity of each liability 2018 and 2019. (If no entry is required for a transaction/event, select "No journal entry required" in the first account field. Do not round intermediate calculations. Enter your answers in whole dollars.)
Answer #1
Date Particulars Debit Credit
2018 $ $
Jan,13 "No Journal entry required"
Feb,1 Cash a/c      84,00,000
         To Paris Bank Loan a/c      84,00,000
May.1 Interst Expenses a/c        2,10,000
Paris Bank Loan a/c      84,00,000
          To Cash a/c      86,10,000
Dec,1 Cash a/c 1,70,64,750
Discount Expenses Payable on Loan a/c      12,35,250
          To Paris Bank Loan a/c 1,83,00,000
Dec,31 Discount Expenses a/c        1,37,250
            To Discount Expenses Payable on Loan a/c        1,37,250
2019
Spet,1 Discount Expenses a/c      10,98,000
            To Discount Expenses Payable on Loan a/c      10,98,000
Spet,1 Paris Bank Loan a/c 1,83,00,000
            To Cash a/c 1,83,00,000
May,1
Interest expenses calculation
$8400000*10%*3/12= $210000
Dec,1
Discounted note
Discount Exp. $18300000*9%*9/12=$1235250
Net Cash Amount received $18300000-$1235250=17064750
Dec,31
Discount exp of 1 month $18300000*9%/12=137250
Sept, 1
Discount exp of 8 month $18300000*9%*8/12=1098000